OpenStudy (anonymous):

Sanjay solved the equation below. Which property did he use to determine that is equivalent to ? distributive property of equality subtraction property of equality multiplication property of equality division property of equality

OpenStudy (anonymous):

7(x+6)=42 7x+42=42 7x=0 x=0

OpenStudy (anonymous):

thats the equation

OpenStudy (anonymous):

Second step the distributive property was used 7(x+6) = 7x + 42 For the third line, 42 was subtracted from both sides, so thats which property?

OpenStudy (anonymous):

A B C OR D?

OpenStudy (anonymous):

In solving the equation, a few properties were used, not just one. The first property was distributive property, then another property by subtracting 42 on both sides of the equation

OpenStudy (anonymous):

a b c or d...

OpenStudy (anonymous):

??

OpenStudy (anonymous):

@skullpatrol